The tube in a medical research X-ray unit uses lead (Z=82) as it's target.

1)

Use the Bohr model to estimate the minimum kinetic energy electrons must have to in order for Kα X-rays to just appear in the X-ray spectrum of the tube.

Ee =

2)

Assuming the electrons are accelerated from rest through a voltage, V, what is the minimum voltage necessary to produce the Kα X-rays?
Answer #1

1) 89229.6 eV

2) 89229.6 Volts

En= -13.6 (2+)2 SE= 13.6 (2-122 cl n2 Kemin = 13.6 (82-1)2 Vol 89229.6 ev = 189.23 ker 189 229.6ev) se = 89.23 x 103 ev 1,6x1